OpenStudy (anonymous):

what is the value of this expression? Do not use negative exponent in your answer

OpenStudy (anonymous):

4^2 * 4^4/ 4^10

OpenStudy (anonymous):

4^2 = 2^4 = 16 4^4 = 2^8 = 256 4^10 = 2^20 = 1048576 Combining things .. \[ 16 \times \frac{256}{1048576} = \frac{1}{256} \]

OpenStudy (anonymous):

thanks so much